A farmer wants to fence in three sides of a rectangular field with 1000 feet of fencing. The other side of the rectangle is a river. If the enclosed area is to be maximum, find the dimensions of the field.
Now I was able to get the area, if I am right, I think the area is 125,000 feet. How do I get the dimensions?

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

A farmer wants to fence in three sides of a rectangular field with 1000 feet of fencing. The other side of the rectangle is a river. If the enclosed area is to be maximum, find the dimensions of the field.
Sides perpendicular to the river:: x ft
Side parallel to the river:: 1000-2x
---------------------
Area of the rectangle:
A = x(1000-2x) = 1000x - 2x^2 sq ft
------
To maximize Area::
A' = -4x + 1000
---
Solve:: -4x + 1000 = 0
x = 250 ft (length of side perpendicular to the river)
1000-2x = 500 ft. (length of the side parallel to the river)
--------------------
Area = 250*500 = 125,000 sq ft
------------------------
Cheers,
Stan H.
---------------------



Answer by Theo(13342)

L = length
W = width

Perimeter is normally equal to 2L + 2W, but you are missing 1 side.
We'll assume it's the length.

since you have 1000 feet of fencing, then L + 2W = 1000

solve for L to get:

L = 1000 - 2W

area is equal to L*W

replace L with 1000 - 2W to get:

area = (1000 - 2W) * W

simplify this to get:

area = 1000W - 2W^2

this is a quadratic equation, so convert it to standard form to get:

area = -2W^2 + 1000W

replace W with x to get:

area = -2x^2 + 1000x

since it's in standard form of y = ax^2 + bx + c:

a = -2
b = 1000
c = 0

since the coefficient of the x^2 term is negative, this will have a maximum.

the maximum point will be at x = -b/(2a) = -1000 / -4 = 250.

the equation will be at a maximum when x = 250

when x = 250, -2x^2 + 1000x = -2(250)^2 + 1000(250) = 125000.

since x represents W which represents the width, then W = 250.

since area is L * W, then L * 250 = 125000.

solve for L to get L = 125000 / 250 = 500

the length is 500 and the width is 250.

the area is 500 * 250 = 125000

the perimeter for fencing is L + 2W = 500 + 2(250) = 500 + 500 = 1000

all is good.